# A Quantum Description of Wave Dark Matter

### v1

@@ Abstract
The above discussion underpins the search for wave DM candidates, such as the axion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. It also raises many questions. How accurate is the result in Eq. (1)? What are the leading order corrections to this result? How can it be derived rigorously? Does the field simply oscillate or can it exhibit large fluctuations around its mean value? When $m\sim 10\,\textrm{eV}$ and DM is neither deep in the wave or particle regime, how should it be described? The goal of the present work is to resolve such questions by providing a fundamentally quantum mechanical description of DM from which the wave picture emerges.
@@ I The density matrix of dark matter
The key question is whether Eq. (7) describes the density matrix for the modes of DM, as has been assumed elsewhere in the literature (see e.g. Ref. [11]). The result holds in many situations within electromagnetism. For example, if a system has thermalized, it has a density matrix in the Gaussian form, although this is not a necessary condition; Eq. (7) is more general. Heuristically, we could imagine that the formation mechanism for the local DM is complex enough – involving virialization and galaxy mergers – that the central limit theorem should apply, and it definitely applies if the system has thermalized. Yet ultralight DM is not expected to have a thermal origin and given the weakness of its interactions, DM’s density matrix may not be in the Gaussian form. We return to consider this question in detail in Sec. VII, in particular we discuss various alternatives for $\hat{\rho}$ and demonstrate how the density matrix can imprint itself in experimental measurements, thereby opening a path to directly accessing it after the discovery of DM. In particular, if DM were in a pure coherent state, $P(\alpha)=\delta(\alpha-\beta)$, its behavior differs considerably from that of the Gaussian distribution in a manner experiments could test. Until then, we generally assume Eq. (7) holds for each mode of the DM field and explore the consequences. We note, however, the full Gaussian form is generally only required to determine the statistical fluctuations of observables, their mean value is often fixed with the weaker condition that $P(\alpha)=P(|\alpha|)$ (this is true of many of the coherence properties we study in Secs. III, IV, and V). We are explicit below regarding which results require a specific $\hat{\rho}$ and which are more general.
@@ II Wave Dark matter as a classical random field
The result is schematic because $\langle|\alpha_{\mathbf{k}}|\rangle\simeq\sqrt{N_{\mathbf{k}}}$ represents only the average behavior; in general we must account for the full statistics of $\alpha_{\mathbf{k}}$ as determined from $P(\alpha_{\mathbf{k}})$. Nevertheless, we note this form is similar to existing efforts to determine the statistical properties of wave DM, in particular see Refs. [28, 35, 29, 36, 37, 38, 39]. The only difference is the exact pattern of statistical fluctuations, which we return to shortly. Pushing further, if we imagine the system is non-relativistic and sufficiently dominated by the single mode of its mass $\omega_{\mathbf{k}}=m$, then we can take $\omega_{\mathbf{k}}\simeq m$, $N/V=\rho/m$, so that at $\mathbf{x}=0$ we have
@@
which satisfy $[\hat{X}_{\mathbf{k}},\hat{Y}_{\mathbf{k}}]=i$ and $\hat{X}_{\mathbf{k}}^{2}+\hat{Y}_{\mathbf{k}}^{2}=2\hat{N}_{\mathbf{k}}+1$. When the classical wave picture applies we can take $\hat{X}_{\mathbf{k}}\simeq\sqrt{2}\operatorname{Re}[\alpha_{\mathbf{k}}]$ and $\hat{Y}_{\mathbf{k}}\simeq\sqrt{2}\operatorname{Im}[\alpha_{\mathbf{k}}]$. The two quadratures are therefore independent Gaussian random fields, and from Eq. (22) directly related to the individual modes of $\phi$ and $\partial\phi$, respectively. As each quadrature of the DM field is normally distributed, if one isolates a single quadrature the statistical properties must differ from a measurement sensitive to both. A further consequence of the quadratures being Gaussian distributed is that number of states in a given mode, $\hat{N}_{\mathbf{k}}\simeq\tfrac{1}{2}(\hat{X}_{\mathbf{k}}^{2}+\hat{Y}_{\mathbf{k}}^{2})$, is distributed as a $\chi^{2}$ distribution with two degrees of freedom; equivalently, it is exponentially distributed. This should be compared to the expectation for particle DM, where $N_{\mathbf{k}}$ is Poisson distributed, corresponding to simply counting the number of particles with a given $\mathbf{k}$. The exponential distribution has a larger variance than Poisson, which can be heuristically associated with constructive and destructive interference of the wave. We establish that these two limits are smoothly connected in Sec. VI.
@@ III Autocorrelation and the Frequency Domain
Figure 1: The DM spatial (left) and temporal (right) autocorrelation functions as defined in Eq. (31), assuming the standard halo model (SHM) as in Eq. (36) with the displayed parameters. In addition to being oscillatory, both functions exhibit a decay that characterizes the coherence of the field as discussed in Sec. IV. On the right we demonstrate that the envelope for the decay of the correlations is accurately represented by Eq. (38) (dashed red), although a simple exponential provides rough approximation (dotted black).
@@
where $f(\mathbf{v})$ and $f(v)$ are the DM velocity and speed distributions, respectively. As a simple ansatz for these distributions, we can adopt the commonly assumed standard halo model (SHM)
@@ IV Coherence of the classical field
Equation (50) lays bare that the coherence time is determined entirely by the energy distribution. In particular, it is sensitive to the inverse width of $p(\omega)$ and diverges as the field becomes dominated by a single frequency. [^16] Of course, in principle $p(\omega)$ could be arbitrarily complicated, such that in the most general case $\tau_{c}$ may only represent a crude measure of the evolution of the autocorrelation function. Nonetheless, if the energy distribution is relative simple such that it is well characterized by its mean and standard deviation (as is the case for DM drawn from the SHM), then the coherence time represents an accurate measure as shown in Fig. 1. To quantify these statements, consider a particularly simple scenario where $p(\omega)$ is a top-hat distribution with mean frequency $\bar{\omega}$ and width $\delta\omega$. The coherence time is then,
@@
Numerically, the size of the correction on the second line is $\simeq 4\times 10^{-7}$ and therefore completely irrelevant. We have included the higher order terms to emphasize that in the approach we have outlined the coherence time is a rigorously defined quantity. Numerically,
@@
Beyond quantitative numerical results, the explicit expressions also allow for the study of the coherence time’s limiting behavior. For instance, based on arguments similar to those in Ref. [29], one may expect the DM coherence time should take the form $\tau_{{\scriptscriptstyle\textrm{DM}}}\sim 1/mv_{0}v_{\odot}$, which is similar to Eq. (53). (We review those arguments at the end of this section.) Nevertheless, the qualitative expression diverges as $v_{\odot}\to 0$, suggesting that an instrument at rest in the halo frame may have an enhanced sensitivity to wave DM. This, of course, cannot be correct. From Eq. (36), when $v_{\odot}\to 0$, the speed distribution has a finite width and therefore should have a finite coherence. Equation (53) manifests our expectation, as we have a finite coherence time of $\tau_{{\scriptscriptstyle\textrm{DM}}}\simeq 4/mv_{0}^{2}$ as the observer’s speed vanishes. For $v_{0}\to 0$, however, the width of the distribution vanishes and the coherence time should diverge, consistent with Eq. (53).
@@ VI Description at the wave-particle boundary
Written in this form, we can immediately read off the probability of observing $k$ quanta,
@@ VII Non-Gaussianities and other forms of the density matrix
So far we have primarily focused on the implications of DM being described by the density matrix with a Gaussian $P(\alpha)$ given in Eq. (7). There are a number of reasons to find this form attractive. Regardless of the state DM was born in the system has undergone considerable evolution. Most significantly, the process of virialization is a violent one. Even if $P(\alpha)$ for DM was non-Gaussian before the galaxy formed, through formation the DM field could have been fragmented and randomized, in which case the local DM field at the present epoch could be treated as a large sum of independent fields, suggesting the quantum central limit theorem could apply. Further, if the DM ever thermalized, then as we review in App. G, there are examples where the evolution to a Gaussian density matrix can be explicitly computed.
Beyond the above suggestive arguments, we offer no proof that the DM density matrix takes a Gaussian form. Resolving this represents an important step in determining the behavior of wave DM and in confirming – or refuting – our various claims. A path to doing so would be study the cosmological evolution of the density matrix for various assumptions of its initial form. There is a wide literature on the topic, delving into the theories of open quantum systems, master equations, and gravitational decoherence [59, 60, 61, 62, 63, 64, 65].
@@ VIII Discussion
Secondly, it will be important to formally establish what the density matrix of DM actually is. We have argued that $\hat{\rho}$ likely takes a Gaussian form and have further noted in Sec. VII that deviations from this would imprint themselves as non-Gaussian fluctuations in DM measurements. Nevertheless, formally one could look to evolve the density matrix of DM from the early Universe until today, which would resolve whether there remains any hint of the state that DM was born in. (In App. G we discuss how the evolution proceeds in a simple example.) Such an analysis could further prove useful in understanding the evolution of DM into configurations such as a BEC.
@@ Appendix A The Quantum Central Limit Theorem
$\tilde{P}(\lambda)\simeq\exp\left[-\frac{1}{4{n}}|\lambda|^{2}\langle|\alpha|^{2}\rangle\right]\!,$ (97)
@@ Appendix G An Example of the Evolution of $P(\alpha)$
As emphasized in the main text, a full justification of the assumed Gaussian $P(\alpha)$ requires understanding how the density matrix of DM evolved. Here we review an example of how one can evolve $P(\alpha)$ in an idealized scenario, one where the DM is coupled to a thermal reservoir. Of course, in this case the fact that we find $P(\alpha)$ tending to a Gaussian is unsurprising: the system is thermalizing and a thermal distribution is equivalent to a Gaussian $P(\alpha)$. Our review follows closely Refs. [84, 64] and we refer to those works for further details. (See also Ref. [59, 60, 61] for a similar study.)

### v2

@@ Abstract
The above discussion underpins the search for wave DM candidates, such as the axion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. It also raises many questions. How accurate is the result in Eq. (1)? What are the leading order corrections to this result? How can it be derived rigorously? Does the field simply oscillate or can it exhibit large fluctuations around its mean value? When $m\sim 10\,\textrm{eV}$ and DM is neither deep in the wave or particle regime, how should it be described? The goal of the present work is to resolve such questions by providing a fundamentally quantum mechanical description of DM from which the wave picture emerges. [^1] Detailed studies of wave DM often employ more sophisticated models than presented in Eq. (1). For instance, Ref. [12] introduced a model to accommodate the finite dispersion of DM by treating the field as a sum of cosines with frequencies weighted by the underlying energy distribution of DM. Nevertheless, we emphasize that even such refinement do not answer the fundamental questions raised.
@@ I The density matrix of dark matter
The key question is whether Eq. (7) describes the density matrix for the modes of DM, as has been assumed elsewhere in the literature (see e.g. Ref. [13]). The result holds in many situations within electromagnetism. For example, if a system has thermalized, it has a density matrix in the Gaussian form, although this is not a necessary condition; Eq. (7) is more general. Heuristically, we could imagine that processes such as galaxy mergers and the violent relaxation associated with virialization [29] are sufficient to render the local DM density a sum of independent distributions to which the central limit theorem applies. Further, if the system ever thermalized during its evolution then the Gaussian form must hold. Yet ultralight DM is not expected to have a thermal origin and given the weakness of its interactions, DM’s density matrix may not be Gaussian. We return to consider this question in detail in Sec. VII, in particular we discuss various alternatives for $\hat{\rho}$ and demonstrate how the density matrix can imprint itself in experimental measurements, thereby opening a path to directly accessing it after the discovery of DM. In particular, if DM were in a pure coherent state, $P(\alpha)=\delta(\alpha-\beta)$, its behavior differs considerably from that of the Gaussian distribution in a manner experiments could test. Until then, we generally assume Eq. (7) holds for each mode of the DM field and explore the consequences. We note, however, the full Gaussian form is generally only required to determine the statistical fluctuations of observables, their mean value is often fixed with the weaker condition that $P(\alpha)=P(|\alpha|)$ (this is true of many of the coherence properties we study in Secs. III, IV, and V). We are explicit below regarding which results require a specific $\hat{\rho}$ and which are more general.
@@ II Wave Dark matter as a classical random field
The result is schematic because $\langle|\alpha_{\mathbf{k}}|\rangle\simeq\sqrt{N_{\mathbf{k}}}$ represents only the average behavior; in general we must account for the full statistics of $\alpha_{\mathbf{k}}$ as determined from $P(\alpha_{\mathbf{k}})$. Nevertheless, we note this form is similar to existing efforts to determine the statistical properties of wave DM, in particular see Refs. [12, 37, 31, 38, 39, 40, 41, 42]. The only difference is the exact pattern of statistical fluctuations, which we return to shortly. Pushing further, if we imagine the system is non-relativistic and sufficiently dominated by the single mode of its mass $\omega_{\mathbf{k}}=m$, then we can take $\omega_{\mathbf{k}}\simeq m$, $N/V=\rho/m$, so that at $\mathbf{x}=0$ we have
@@
which satisfy $[\hat{X}_{\mathbf{k}},\hat{Y}_{\mathbf{k}}]=i$ and $\hat{X}_{\mathbf{k}}^{2}+\hat{Y}_{\mathbf{k}}^{2}=2\hat{N}_{\mathbf{k}}+1$. When the classical wave picture applies we can take $\hat{X}_{\mathbf{k}}\simeq\sqrt{2}\operatorname{Re}[\alpha_{\mathbf{k}}]$ and $\hat{Y}_{\mathbf{k}}\simeq\sqrt{2}\operatorname{Im}[\alpha_{\mathbf{k}}]$. The two quadratures are therefore independent Gaussian random fields, and from Eq. (22) directly related to the individual modes of $\phi$ and $\partial\phi$, respectively. As each quadrature of the DM field is normally distributed, if one isolates a single quadrature the statistical properties must differ from a measurement sensitive to both. A further consequence of the quadratures being Gaussian distributed is that the number of states in a given mode, $\hat{N}_{\mathbf{k}}\simeq\tfrac{1}{2}(\hat{X}_{\mathbf{k}}^{2}+\hat{Y}_{\mathbf{k}}^{2})$, is distributed as a $\chi^{2}$ distribution with two degrees of freedom; equivalently, it is exponentially distributed. This should be compared to the expectation for particle DM, where $N_{\mathbf{k}}$ is Poisson distributed, corresponding to simply counting the number of particles with a given $\mathbf{k}$. The exponential distribution has a larger variance than Poisson, which can be heuristically associated with constructive and destructive interference of the wave. We establish that these two limits are smoothly connected in Sec. VI.
@@ III Autocorrelation and the Frequency Domain
Figure 1: The DM spatial (left) and temporal (right) autocorrelation functions as defined in Eq. (31), assuming the standard halo model (SHM) as in Eq. (36) with the displayed parameters. In addition to being oscillatory, both functions exhibit a decay that characterizes the coherence of the field as discussed in Sec. IV. On the right we demonstrate that the envelope for the decay of the correlations is accurately represented by Eq. (38) (dashed red), although a simple exponential provides a rough approximation (dotted black).
@@
where $f(\mathbf{v})$ and $f(v)$ are the DM velocity and speed distributions, respectively, and we integrate over all values of $\mathbf{v}$ and $v\geq 0$. As a simple ansatz for these distributions, we can adopt the commonly assumed standard halo model (SHM)
@@ IV Coherence of the classical field
Equation (50) lays bare that the coherence time is determined entirely by the energy distribution. In particular, it is sensitive to the inverse width of $p(\omega)$ and diverges as the field becomes dominated by a single frequency. [^17] Of course, in principle $p(\omega)$ could be arbitrarily complicated, such that in the most general case $\tau_{c}$ may only represent a crude measure of the evolution of the autocorrelation function. Nonetheless, if the energy distribution is relative simple such that it is well characterized by its mean and standard deviation (as is the case for DM drawn from the SHM), then the coherence time represents an accurate measure as shown in Fig. 1. To quantify these statements, consider the particularly simple scenario of a top-hat distribution – $p(\omega)=(1/\delta\omega)\Theta(\omega-\bar{\omega}+\delta\omega/2)\Theta(\bar{\omega}+\delta\omega/2-\omega)$ – with mean frequency $\bar{\omega}$ and width $\delta\omega$. The coherence time is then,
@@
Numerically, the second line is $\simeq 1+4\times 10^{-7}$ and therefore even the ${\cal O}(v^{2})$ corrections are completely irrelevant. We have included the higher order terms to emphasize that in the approach we have outlined the coherence time is a rigorously defined quantity. Numerically,
@@
Beyond quantitative numerical results, the explicit expressions also allow for the study of the coherence time’s limiting behavior. For instance, based on arguments similar to those in Ref. [31], one may expect the DM coherence time should take the form $\tau_{{\scriptscriptstyle\textrm{DM}}}\sim 1/mv_{0}v_{\odot}$, which is similar to Eq. (53). (We review those arguments at the end of this section.) Nevertheless, the qualitative expression diverges as $v_{\odot}\to 0$, suggesting that an instrument at rest in the halo frame may have an enhanced sensitivity to wave DM. This, of course, cannot be correct and our exact expressions reveal the appropriate behavior. From Eq. (36), when $v_{\odot}\to 0$, the speed distribution has a finite width and therefore should have a finite coherence. Equation (53) manifests our expectation, as we have a finite coherence time of $\tau_{{\scriptscriptstyle\textrm{DM}}}\simeq 4/mv_{0}^{2}$ as the observer’s speed vanishes. For $v_{0}\to 0$, however, the width of the distribution vanishes and the coherence time should diverge, consistent with Eq. (53).
@@ V Higher Order Coherence
where as noted above we integrate over $\tau\in(-\infty,\infty)$.
@@ VI Description at the wave-particle boundary
Importantly, the density matrix is diagonal in the Fock basis. We emphasize this is not generic and it is certainly not true for a pure coherent state, see App. G for further discussion. Regardless, it implies that we can immediately read off the probability of observing $k$ quanta,
@@ VII Non-Gaussianities and other forms of the density matrix
So far we have primarily focused on the implications of DM being described by the density matrix with a Gaussian $P(\alpha)$ given in Eq. (7). There are a number of reasons to find this form attractive. Regardless of the state DM was born in the system has undergone considerable evolution. Most significantly, the process of virialization is a violent one [29]. Even if $P(\alpha)$ for DM was non-Gaussian before the galaxy formed, through formation the DM field could have been fragmented and randomized, in which case the local DM field at the present epoch could be treated as a large sum of independent fields, suggesting the quantum central limit theorem could apply. Further, if the DM ever thermalized, then as we review in App. G, there are examples where the evolution to a Gaussian density matrix can be explicitly computed.
Beyond the above suggestive arguments, we offer no proof that the DM density matrix takes a Gaussian form. Resolving this represents an important step in determining the behavior of wave DM and in confirming – or refuting – our various claims. A path to doing so would be to study the cosmological evolution of the density matrix for various assumptions of its initial form. There is a wide literature on the topic, delving into the theories of open quantum systems, master equations, and gravitational decoherence [62, 63, 64, 65, 66, 67, 68].
@@ VIII Discussion
Secondly, it will be important to formally establish what the density matrix of DM actually is. We have argued that $\hat{\rho}$ likely takes a Gaussian form and have further noted in Sec. VII that deviations from this would imprint themselves as non-Gaussian fluctuations in DM measurements. Nevertheless, formally one could look to evolve the density matrix of DM from the early Universe until today, which would resolve whether there remains any hint of the state that DM was born in. (In App. G we discuss how the evolution proceeds in a particularly simple example.) Such an analysis could further prove useful in understanding the evolution of DM into configurations such as a BEC.
@@ Appendix A The Quantum Central Limit Theorem
$\tilde{P}(\lambda)\simeq\exp\left[-\frac{1}{4}|\lambda|^{2}\langle|\alpha|^{2}\rangle\right]\!,$ (97)
@@ Appendix G An Example of the Evolution of $P(\alpha)$
As emphasized in the main text, a full justification of the assumed Gaussian $P(\alpha)$ requires understanding how the density matrix of DM evolved. Here we review an example of how one can evolve $P(\alpha)$ in a highly idealized scenario, one where the DM is coupled to a thermal reservoir. We emphasize that in no sense is this intended to be a model for how the $P(\alpha)$ of DM evolved in our Universe and the final result is unsurprising: $P(\alpha)$ evolves to a Gaussian as the system is thermalizing and a thermal distribution is equivalent to a Gaussian $P(\alpha)$. Our review follows closely Refs. [87, 67] and we refer to those works for further details. (See also Ref. [62, 63, 64] for a similar study.)
@@
Lastly, we can use the above example as an opportunity to expand on the point that the Gaussian $P(\alpha)$ corresponds to a density matrix that is diagonal in the Fock basis, as seen in Eq. (72). This is far from generic. A pure coherent state $|\alpha_{0}\rangle\langle\alpha_{0}|$ has support for arbitrary off-diagonal terms,
$\hat{\rho}=e^{-|\alpha_{0}|^{2}}\sum_{n,k=0}^{\infty}\frac{\alpha_{0}^{n}(\alpha_{0}^{*})^{k}}{\sqrt{n!k!}}|n\rangle\langle k|.$ (149)
As another way to view this, neglecting interactions, the number operator commutes with the Hamiltonian and therefore the Fock basis can also be viewed as the energy basis: the Gaussian is diagonal in this space whilst generic states are not. Using Eq. (148) we can explicitly compute the evolution of both the diagonal and off-diagonal contributions to Eq. (149) as the system thermalizes, which can be done analytically in terms of sums of Hermite polynomials. Doing so one finds that once $t\sim\gamma^{-1}$ the off-diagonal terms begin to be suppressed leaving only the Gaussian supported diagonal elements. Although we have only studied this in a simple model, the diagonalization in energy will be a generic feature of the evolution of systems towards a Gaussian $P(\alpha)$.